# The heat flow on glued spaces with varying dimension

### v1

@@ Abstract
We examine under which conditions the canonical heat flow on glued manifolds is ergodic and irreducible. Glued manifolds are spaces consisting of manifolds of varying dimension connected by a weakly doubling measure. Moreover, we construct a non-local perimeter functional, the heat excess, to raise the question of its $\Gamma-$convergence to the standard perimeter functional.
@@ 1 Motivation
The heat equation is one of the most important and most often applied partial differential equations. This equation has applications in a vast amount of areas and fields. In this note, we analyze the existence of an ergodic heat flow on so called weighted glued manifolds. These include spaces which consist of manifolds with different dimensions that are connected through weights. The idea behind these spaces is to incorporate a general setting for e.g. manifold learning and similar applications where data is distributed locally in the shape of manifolds but can vary globally. The weights make sure that each component distributes in the same way and that the data is overall continuous in some sense. To state our main theorem, Theorem 2.2, in an informal way, we will prove
@@
Let $(X,d,\mu)$ be the union of weighted smooth manifolds with transversal intersections s.t. $\mu$ fulfills a weak doubling condition. Then, the canonical heat flow on $X$ is ergodic and irreducible.
Since the heat flow is ergodic and irreducible it is possible to investigate the short-term behavior of the heat flow and investigate the $\Gamma-$ (or Mosco-)convergence of a non-local perimeter functional to the local perimeter. With this in hand, one can construct a classifier on these spaces. Moreover, the definitions carry over to a more general setting of metric measure spaces instead of smooth compact manifolds. For a more details, see 5.
1.1 Structure of the note
@@
Additionally, we recapitulate the standard theorems and properties which we apply later. The advanced theorems and notations are introduced in their respective sections as to not overload Section 2.
@@
Roughly, the idea is to use the local structure of the manifolds to show that only constants remain static under the flow. Together with spectral theory of the resolvent associated with the flow, this allows to prove a spectral gap and exponential convergence to constants for every initial condition. This convergence, in turn, implies ergodicity and irreducibility.
@@
Next, in Section 5, we present the definitions and basic theory needed to state a conjecture about the $\Gamma-$convergence (or Mosco-convergence) of a non-local perimeter functional to the perimeter functional in glued spaces and general metric measure spaces. This functional is known as the heat excess and uses the heat flow constructed beforehand.
Finally, in Section 6, we discuss related work in the different areas of this note and connections to the different fields of study where these problems occur.
@@ 2 Setting and main theorem
We will now give the setting for the metric measure space $(X,d,\mu)$.
Our domain $X$ is the union of two embedded Euclidean manifolds that can vary in dimension and intersect. To be precise,
$X:=M_{1}\cup M_{2}$
where $M_{i}$ are compact, connected, oriented manifolds with smooth boundary. The dimension of $M_{i}$ is $n_{i}$ and the intersection $\emptyset\neq L:=M_{1}\cap M_{2}$ lies in the inner of $M_{i}$ and is a smooth manifold of dimension $k$. This is the case if e.g. the intersection is transversal in the surrounding space or a smooth manifold containing an open neighborhood of the intersection, c.f. [29, Chapter 5].
The (intrinsic) distance $d$ is defined as the glued distance which is obtained by taking the infimum over the length of all connecting paths between points. All balls $B$ in this note are w.r.t. the intrinsic distance $d$ on $X$.
Lastly, $\mu$ takes the role of a volume measure on $X$ that connects the manifolds and is composed of the weighted volume measures
$\mu:=\omega_{1}\cdot\text{vol}_{M_{1}}+\omega_{2}\cdot\text{vol}_{M_{2}}$
where the weights $\omega_{i}\in A_{2}(M_{i}),\omega_{i}>0$ are Muckenhoupt, see Definition 2.9. Additionally, we require that $\mu(L)=0$.
@@
This ensures a weak type of continuity of $\mu$.
@@
The $N$-doubling condition can be formulated with the Euclidean (extrinsic) distance. In this case one would need to strengthen the transversality into a metric-quantitative version, i.e. there exist balls near the intersection which do not intersect other manifolds.
@@
Alternatively, one could consider another connection of the manifolds instead of the weights. It is possible to get a continuity over the intersection by collapsing an area around it into a lower dimensional submanifold and changing the distances and topology accordingly. In this note, we decided for the approach with weights as it satisfies a continuity over the whole domain and fits more naturally into the framework of manifold learning.
Now, we are able to state the main theorem of this note.
@@
Let $X=M_{1}\cup M_{2}$ with the intrinsic distance $d$ be a given transversal union of two smooth compact manifolds with boundary s.t. the intersection $L:=M_{1}\cap M_{2}$ lies in the inner of both of them. Moreover, assume non-negative weights $\omega_{1}\in A_{2}(M_{1})$ and $\omega_{2}\in A_{2}(M_{2}),\omega_{i}>0$ s.t. $\omega_{2}$ is bounded from above and below with $k=n_{2}-1$ where w.l.o.g. $n_{1}\geq n_{2}$. Additionally, assume that
$\mu:=\omega_{1}\cdot\text{vol}_{M_{1}}+\omega_{2}\cdot\text{vol}_{M_{2}}$
@@
The construction of the heat flow is done via its semigroup which itself stems from the natural Dirichlet form $\mathcal{E}$ on $X$:
$\operatorname*{\mathcal{E}}(f):=\int\limits_{X}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu:=\int\limits_{M_{1}}|\nabla^{M_{1}}f|^{2}\mathop{}\!\mathrm{d}\mu_{1}+\int\limits_{M_{2}}|\nabla^{M_{2}}f|^{2}\mathop{}\!\mathrm{d}\mu_{2}.$
This is to say that $\operatorname*{\mathcal{E}}$ is the manifold gradient on each manifold $M_{i}$ combined with the partial measure $\mu_{i}:=\omega_{i}\cdot\text{vol}_{M_{i}}$. For a general definition of Dirichlet forms and their properties see Definition 2.4 and following.
@@
Figure 1: Example of a glued manifold
Example 2.3.
A typical example would be a one-dimensional line intersecting a two-dimensional disk with the density $\frac{1}{|x|}$ that lifts the dimensions at their intersection. This is so that their dimensions don’t differ too much and the measure is continuous in some way. A visualization can be seen in Figure 1. We will see that if the difference in dimension at their intersection is less than one, the heat flow connects both of them and is ergodic. In the case that the difference in their dimension is too large, the capacity of the intersection w.r.t. one of both manifolds is zero and thus the heat flow does not need to be ergodic.
Now, we will give the precise definition of Dirichlet forms as well as of Muckenhoupt weights. The following definitions can be compared to [20, 60, 62].
Definition 2.4 (Dirichlet form).
A symmetric form (non-negative definite densely defined symmetric bilinear form) $(\mathcal{E},D(\mathcal{E}))$ on $L^{2}$ is called a Dirichlet form if it is closed and Markovian, i.e.
$D(\mathcal{E})\text{ with norm }\mathcal{E}_{1}(u):=\left\lVert u\right\rVert_{L^{2}}+\mathcal{E}(u,u)\text{ is complete}$
@@
A core $\mathcal{C}$ of $\mathcal{E}$ is a dense subset of $D(\mathcal{E})\cap C^{0}$ (in $D(\mathcal{E})$ w.r.t. $\mathcal{E}_{1}$ and in $C^{0}$ w.r.t. $L^{\infty}-$norm) and we call $\mathcal{E}$ regular if a core exists.
Lastly, a Dirichlet form $\mathcal{E}$ is called strongly local if for $u,v\in D(\mathcal{E})$ compactly supported and $u$ constant in a neighborhood of $v$: $\mathcal{E}(u,v)=0$. Furthermore, it is called strictly local if the associated energy metric $d_{\mathcal{E}}$ (see below) is a metric and induces the same topology as $d$.
We will write $\mathcal{E}(u)$ for $\mathcal{E}(u,u)$. For closed forms, the Markovian condition can be rephrased as $\mathcal{E}$ being non-increasing under normal contractions.
Now, we will define the energy metric and what we mean when we call a function (locally) Lipschitz. On the core and thus by extension on the domain, we define the energy measure $\Gamma(u)$ which is a Radon measure, see [60, 62], via:
$\int_{X}\varphi\mathop{}\!\mathrm{d}\Gamma(u):=\mathcal{E}(\varphi u,u)-\frac{1}{2}\mathcal{E}(u^{2},\varphi).$
This can also be seen as a carré-du-champ-operator if ones looks at the bilinear form obtained using a polarization formula and for $u,v\in D(\mathcal{E})$
$\int_{X}\varphi\mathop{}\!\mathrm{d}\Gamma(u,v)=\frac{1}{2}\left(\mathcal{E}(\varphi u,v)+\mathcal{E}(\varphi v,u)-\frac{1}{2}\mathcal{E}(\varphi,uv)\right).$
@@
$d_{\mathcal{E}}(x,y):=\sup\left\{u(x)-u(y)\left|u\in\mathcal{C},\frac{\mathop{}\!\mathrm{d}\Gamma(u)}{\mathop{}\!\mathrm{d}\mu}\leq 1\right.\right\}.$
This notion means that $\Gamma(u)\ll\mu$ and that the Radon-Nikodym derivative is bounded by $1$. A priori, this energy metric does not have to be a metric as it can be infinite or zero for different $x,y$.
We note that $\mathcal{E}$ in the form above for the setting of manifolds is in fact a Dirichlet form and is regular, strongly local and strictly local. In particular, we could also use the energetic metric instead of the glued distance.
Finally, we call a function $u\in D(\mathcal{E})$ Lipschitz, $u\in\operatorname*{Lip}$, if $\Gamma(u)\ll\mu$ and the Radon-Nikodym derivative is bounded. This bound is then called the Lipschitz constant. In this case, we write $|\nabla u|^{2}:=\frac{\mathop{}\!\mathrm{d}\Gamma(u)}{\mathop{}\!\mathrm{d}\mu}$ and $|\nabla u|:=\sqrt{|\nabla u|^{2}}$. This coincides with our intuition that $\mathcal{E}$ is the $H^{1}-$seminorm and $\Gamma$ the square of the derivative.
For the connectedness of $(X,d,\mu)$ we will need a Poincare inequality.
Definition 2.5 ($2$-Poincare inequality).
We say that $X$ satisfies a $2$-Poincare inequality if for each $f\in\operatorname*{Lip}$ there exists $C>0$ and a dilation factor $\lambda\geq 1$ s.t. for each ball $B\subset X$:
$\fint_{B}|f-f_{B}|^{2}\mathop{}\!\mathrm{d}\mu$ $\leq Cr^{2}\fint_{\lambda B}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu.$
@@
Note, if $X$ satisfies a Poincare inequality it has to be connected up to negligible sets. Assume it would not be connected up to a null set and let $X_{1}$ be one component with $f=\chi_{X_{1}}$ and $|\nabla f|=0$. Then, this would be a contradiction as $\mu(X_{1})>0$ and $01>f_{B}>$ 0 for $B$ large enough.
If not specified we will assume the dilation factor $\lambda$ to be one.
The heat flow is defined using its semigroup which originates in the Dirichlet form. In summary, this can be done as follows. Given the Dirichlet form $\operatorname*{\mathcal{E}}$, we define the Laplace-Beltrami operator $\operatorname*{\mathcal{A}}$ with its domain $D(\operatorname*{\mathcal{A}})$ in the following way
$D(\operatorname*{\mathcal{A}}):=\{f\in D(\operatorname*{\mathcal{E}})|\exists\operatorname*{\mathcal{A}}f\in L^{2}:(\operatorname*{\mathcal{A}}f,g)=-\operatorname*{\mathcal{E}}(f,g)\;\forall g\in D(\operatorname*{\mathcal{E}})\}.$
By definition, this is a selfadjoint symmetric non-positive operator which induces an analytic semigroup $S_{h}$ which we call the heat semigroup.
In Section 3, we will also show continuity and ergodicity properties of these objects. For this, we will need capacity bounds.
@@
The relative $2-\mu-$capacity is defined on compact sets and extended to measurable sets in the usual way:
$\text{Cap}_{2,\mu}(K,\Omega):=\inf\left\{\left.\int\limits_{\Omega}|\nabla u|^{2}\mathop{}\!\mathrm{d}\mu\right|u\in\operatorname*{Lip}\cap C_{c}(\Omega),u\geq\mathds{1}_{K}\right\}.$
@@
Additionally, we want to show that our constructed heat flow connects the whole space. The right notion to make this precise is the notion of ergodicity and irreducibility, see [20]. In the following, let $S_{h}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ be a semigroup.
@@
We call a measurable set $E\subset X$ $S_{h}$ invariant if for all $f\in L^{2},h>0$ we have
[…]
Then, $S_{h}$ is irreducible if for all $S_{h}$ invariant sets $E$ we have $\mu(E)=0$ or $\mu(E^{c})=0$.
@@
We call $S_{h}$ ergodic if the set function $T_{t}$ defined by
$T_{t}:X\to X,\quad E\mapsto\operatorname*{supp}(S_{t}\mathds{1}_{E})$
is ergodic for each fixed $t>0$, i.e. if for a measurable set $E$ with $T_{t}^{-1}(E)=E$ we have $\mu(E)=0$ or $\mu(E^{c})=0$.
@@
We say $\omega$ is in the class $A_{2}(M)$ of $2-$Muckenhoupt weights on a smooth compact manifold $M$ if for all balls $B$
@@
For the typical weight $\omega=|x|^{\alpha}$ in $\mathbb{R}^{d}$, we have $\omega\in A_{2}$ iff $\alpha\in(-n,n)$.

### v2

@@ Abstract
In this paper, we introduce a new concept of glued manifolds and investigate under which conditions the canonical heat flow on these glued manifolds is ergodic and irreducible. Glued manifolds are metric spaces consisting of manifolds of varying dimension connected by a weakly doubling measure. This can be seen as a condition on the jump in dimension. From another perspective, this construction also defines the Brownian motion on these glued spaces. Using the heat flow, we construct a nonlocal perimeter functional, the heat excess, to raise the question of its $\Gamma$-convergence to the standard perimeter functional.
@@ 1 Motivation
The heat equation is one of the most important and most often applied partial differential equations. This equation has applications in a large number of areas and fields. In this paper, we analyze the existence of an ergodic heat flow on so-called weighted glued manifolds. These include spaces consisting of manifolds with different dimensions that are connected through weights. The idea behind these spaces is to incorporate a general setting, for instance, for manifold learning and similar applications where data is distributed locally in the shape of manifolds but can vary globally. This assumption is also called the union of manifold hypothesis. To be more precise, the data in machine learning is often given through points in a very high dimensional space. For instance, in the classical example of MNIST, a data set consisting of hand-drawn digits, one has a training data set of 60.000 digits which are represented by $28\times 28$ gray scale images. The canonical embedding space is $\mathbb{R}^{784}$. However, data can be classified by algorithms with comparatively few parameters. This success gives rise to the manifold hypothesis, namely that the data possesses an intrinsic lower dimensional structure. In the example of MNIST, one could think of a few parameters describing the structure of each digit (for $8$ this would be two parameterized loops). Here, one can already see that not all digits have the same variety. For example, a 0 has less features than an $8$. The union of manifold hypothesis suggests that one should think of the data as combined manifolds of different dimensions. This is supported by numerical simulations [15].
The weights in our formulation ensure that each component contributes in the same way and that the data are overall continuous in a weak sense, which is made precise later. To state our main theorem, Theorem 2.5, in an informal way, we prove the following:
@@
Let $(X,d,\mu)$ be the union of weighted smooth manifolds with transversal intersections such that $\mu$ satisfies a weak doubling condition. Then, the canonical heat flow in $X$ is ergodic and irreducible.
The heat flow here is defined through the Dirichlet form given from the manifolds with the combined volume measure. This definition leads to the Laplace-Beltrami operator on both manifolds individually and has a Dirichlet and Neumann compatibility condition at the intersection, that is, the inflow and outflow coincide, which ensures mass preservation. This also constructs the canonical Brownian motion as the process with the Laplace-Beltrami operator as the generator. The ergodicity and irreducibility follows from a spectral gap obtained for the Laplace-Beltrami operator.
The weak doubling condition heuristically says that the jump in dimension between the manifolds cannot be too high. If the effective difference, which is the difference with consideration of the weights, is larger than one, the flow will separate and the manifolds will not be connected.
Since the heat flow is ergodic and irreducible, it is possible to investigate the short-term behavior of the heat flow and ask the question of the $\Gamma$- (or Mosco-)convergence of a nonlocal perimeter functional to the local perimeter. With this in hand, one can construct a classifier on these spaces. Moreover, the definitions carry over to a more general setting of metric measure spaces instead of smooth compact manifolds. For more details, see Section 5.
In Section 6, we compare our results with existing work and present related settings and papers.
1.1 Structure of the paper
@@
Additionally, we recapitulate the standard theorems and properties which we apply later. The advanced theorems and notations are introduced in their respective sections to avoid overloading Section 2.
@@
Basically, the idea is to use the local structure of the manifolds to show that only constant functions remain static under the flow. Together with spectral theory of the resolvent associated with the flow, this allows us to prove a spectral gap and exponential convergence to constants for every initial condition. This convergence, in turn, implies ergodicity and irreducibility.
@@
Next, in Section 5, we present the definitions and basic theory needed to state a conjecture about the $\Gamma$-convergence (or Mosco-convergence) of a nonlocal perimeter functional to the perimeter functional in glued spaces and general metric measure spaces. This functional is known as the heat excess and uses the heat flow constructed beforehand.
Finally, in Section 6, we discuss related work in the different areas of this paper and connections to the different fields of study where these problems occur.
@@ 2 Setting and Main Theorem
We now present the setting for the metric measure space $(X,d,\mu)$. We first state the results for two compact manifolds and then extend to a generalized setting.
Our domain $X$ is the union of two embedded smooth Riemannian manifolds that can vary in dimension and intersect. To be precise,
$X\coloneqq M_{1}\cup M_{2}.$
where $M_{i}$ are compact connected Riemannian manifolds with a smooth boundary. We assume that the dimension of $M_{i}$ is $n_{i}$ and the intersection $\emptyset\neq L\coloneqq M_{1}\cap M_{2}$ lies within $M_{i}$ and that $L$ is a smooth manifold of dimension $k\leq\min(n_{1},n_{2})$. This is the case if, for instance, the intersection is transversal in the surrounding space or a smooth manifold containing an open neighborhood of the intersection, see Definition 2.2.
The distance $d$ is defined as the glued distance defined as follows, where $d_{i}$ is the distance in $M_{i}$.
Definition 2.1 (Glued distance).
For the manifolds $M_{1},M_{2}$, we define the glued distance $d$ as:
$d(x,y)$ $\coloneqq\inf\left\{\sum\limits_{i=0}^{N-1}\widetilde{d}(x_{i},x_{i+1}):x_{0}=x,x_{N}=y\right\},$
$\widetilde{d}(x,y)$ $\coloneqq\begin{cases}\min(d_{1}(x,y),d_{2}(x,y))&\text{for }x,y\in L,\\ d_{1}(x,y)&\text{for }(x,y)\in M_{1}\times M_{1},(x,y)\not\in M_{2}\times M_{2},\\ d_{2}(x,y)&\text{for }(x,y)\not\in M_{1}\times M_{1},(x,y)\in M_{2}\times M_{2},\\ \infty&\text{for }(x,y)\not\in(M_{1}\times M_{1})\cup(M_{2}\times M_{2}).\end{cases}$
which is obtained by taking the infimum over the length of all connecting paths between points. This distance restricted to each manifold $M_{i}$ is equivalent to the original distance on the manifolds, since the intersection $L$ is a compact smooth submanifold. All metric balls $B$ in this chapter are considered with respect to this distance $d$ on $X$ if not stated otherwise. But we show that other metrics induce the same topology and even are equivalent, see the comment after Proposition 2.8.
Lastly, $\mu$ takes the role of a volume measure on $X$ that connects the manifolds and is the sum of the weighted volume measures, namely
$\mu\coloneqq\omega_{1}\cdot\text{vol}_{M_{1}}+\omega_{2}\cdot\text{vol}_{M_{2}}$
where the weights $\omega_{i}\in A_{2}(M_{i})$, $\omega_{i}>0$ are Muckenhoupt, see Definition 2.13. Additionally, we require that $\mu(L)=0$ and that the manifolds $M_{i}$ are transversal. Intuitively, this says that the manifolds do not coincide or go in the same direction in a local neighborhood near the intersection.
Definition 2.2 (Transversality, [34, Section 1.5]).
Two manifolds $M_{1},M_{2}$ are transversal if
  • there exists a smooth manifold $M$ and a local neighborhood $U$ containing the intersection $L$ such that $M_{i}\cap U$ is a smooth submanifold of $M\cap U$ of dimension $n_{i}$
  • for all $x\in L$, we have a relation of the tangent spaces:
    $T_{x}M_{1}+T_{x}M_{2}=T_{x}M.$
Given this definition, there is a direct consequence which states the basic properties of the intersection.
Proposition 2.3 ([34, Section 1.5]).
If $M_{1}$ and $M_{2}$ are transversal in $M$, their intersection $L$ is again a submanifold of $M$ with
$\text{codim}(L)=\text{codim}(M_{1})+\text{codim}(M_{2})$
or equivalently, $k=n_{1}+n_{2}-n$.
@@
This ensures a weak type of continuity of $\mu$. This type of condition is the best we can hope for in this class of spaces, as a doubling condition would imply that the whole space has the same dimension. Moreover, there is a rich class of spaces that fulfill this type of generalized doubling condition. We provide examples in Example 2.6 and Example 4.2.
@@
The $N$-doubling condition can also be formulated with the Euclidean (extrinsic) distance. In this case, one would need to strengthen the transversality into a metric-quantitative version, i.e., there exist balls near the intersection which do not intersect the other manifold.
@@
Alternatively, one could consider another connection of the manifolds instead of the weights. It is possible to get a continuity over the intersection by collapsing an area around it into a lower-dimensional submanifold and changing the distances and topology accordingly. In this paper, we use the approach with weights as it satisfies a continuity over the whole domain and fits more naturally into the framework of manifold learning. With continuity, we mean that the effective dimension (including potentially vanishing or singular weights) does not jump and changes continuously. This ensures capacity bounds and trace conditions as proven in Section 3.
Now, we are able to state the main theorem of this paper.
@@
Let $X=M_{1}\cup M_{2}$ with the glued distance $d$ be a given transversal union of two smooth compact manifolds with boundary such that intersection $\emptyset\neq L\coloneqq M_{1}\cap M_{2}$ lies in the inner of both of them with $\text{dim}(M_{i})=n_{i}$, $\text{dim}(L)=k$. Moreover, assume nonnegative weights $\omega_{1}\in A_{2}(M_{1})$ and $\omega_{2}\in A_{2}(M_{2})$, $\omega_{i}>0$ such that $\omega_{2}$ is bounded from above and below, $\omega_{1}$ is bounded in a neighborhood of $\partial_{1}M_{1}$ and $k=n_{2}-1$ where without loss of generality $n_{1}\geq n_{2}$. Additionally, assume that
$\mu\coloneqq\omega_{1}\cdot\text{vol}_{M_{1}}+\omega_{2}\cdot\text{vol}_{M_{2}}$
@@
This theorem can be generalized in a form that allows for unbounded weights on both manifolds and for multiple manifolds that intersect. This is done in Section 4, Theorem 4.1, and follows the same ideas and steps as the proof of Theorem 2.5. For clarity and to be better readable, we first state the simpler version and prove this.
The construction of the heat flow is done via its semigroup, which itself originates from the natural Dirichlet form $\mathcal{E}$ on $X$:
$\operatorname*{\mathcal{E}}(f)$ $\coloneqq\int\limits_{X}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu\coloneqq\int\limits_{M_{1}}|\nabla^{M_{1}}f|^{2}\mathop{}\!\mathrm{d}\mu_{1}+\int\limits_{M_{2}}|\nabla^{M_{2}}f|^{2}\mathop{}\!\mathrm{d}\mu_{2},$
$D(\operatorname*{\mathcal{E}})$ $\coloneqq\overline{C^{\infty}(\mathbb{R}^{\widetilde{n}})}^{\sqrt{\left\lVert\cdot\right\rVert_{L^{2}(X,\mu)}^{2}+\operatorname*{\mathcal{E}}(\cdot)}}=:H^{1}_{*}.$
Here, $\widetilde{n}$ is the dimension of the surrounding space in which $X$ is embedded.
That $\operatorname*{\mathcal{E}} is$ the manifold gradient in each manifold $M_{i}$ combined with the partial measure $\mu_{i}\coloneqq\omega_{i}\cdot\text{vol}_{M_{i}}$. Sometimes, this Dirichlet form is also called Cheeger energy; see, for example, [16]. We call the individual Dirichlet forms $\operatorname*{\mathcal{E}}_{i}(f)\coloneqq\int_{M_{i}}|\nabla^{M_{i}}f|^{2}\mathop{}\!\mathrm{d}\mu_{i}$. For a general definition of Dirichlet forms and their properties, see Definition 2.7 and the following.
$M_{1}$$M_{2}$$\omega_{1}=\frac{1}{|x|}$
Figure 1: Example of a glued manifold consisting of the weighted manifolds $M_{1}$ and $M_{2}$.
Example 2.6.
A typical example would be a one-dimensional line intersecting a two-dimensional disk with the density $\frac{1}{|x|}$ that lifts the dimensions at their intersection. This is done to ensure that their dimensions do not differ too much and that the measure is continuous in the sense of an $N$-doubling condition.
A visualization can be seen in Figure 1. We show that if the difference in dimension at their intersection is less than one, the heat flow connects both of them and is ergodic. This is an example of a transversal intersection as one can take $\mathbb{R}^{3}$ as the surrounding manifold and compute $\operatorname*{span}\{e_{1},e_{2}\}+\operatorname*{span}\{e_{3}\}=\mathbb{R}^{3}$.
In the case that the difference in their dimension is too large, the capacity of the intersection with respect to one of both manifolds is zero, and thus the heat flow does not need to be ergodic. For this precise example, the disk and line each without density, the flow would be separate on the disk and the line. This is due to the fact that the intersection is a point and thus has capacity zero in the disk, i.e., it is too small for the flow and the Brownian motion meets it with probability zero, see [58, Theorem 3.2]. Similarly, from the perspective of the line, the mass of the disk near the intersection is not enough.
Now, we give the precise definition of Dirichlet forms as well as of Muckenhoupt weights. The following definitions can be compared to [24, 64, 68].
Definition 2.7 (Dirichlet form).
A symmetric form (nonnegative definite densely defined symmetric bilinear form) $(\mathcal{E},D(\mathcal{E}))$ on $L^{2}$ is called a Dirichlet form if it is closed and Markovian, i.e.,
@@
A core $\mathcal{C}$ of $\mathcal{E}$ is a dense subset of $D(\mathcal{E})\cap C^{0}$ in the sense that $\mathcal{C}$ is dense in the domain $D(\mathcal{E})$ with respect to $\left\lVert u\right\rVert_{H^{1}_{*}}$ and in $C^{0}$ with respect to $L^{\infty}$-norm. We call $\mathcal{E}$ regular if a core exists.
Lastly, a Dirichlet form $\mathcal{E}$ is called strongly local if for $u,v\in D(\mathcal{E})$ compactly supported and $u$ constant in a neighborhood of $v$: $\mathcal{E}(u,v)=0$. Furthermore, it is called strictly local (sometimes also called strongly regular) if the associated energy metric $d_{\mathcal{E}}$ (see below) is a metric and induces the same topology as $d$.
Remark.
We give a reason for the notation $\left\lVert u\right\rVert_{H^{1}_{*}}$ in Section 3.
We write $\mathcal{E}(u)$ for $\mathcal{E}(u,u)$. For closed forms, the Markovian condition can be rephrased as $\mathcal{E}$ being nonincreasing under all normal contractions.
Now, we define the energy metric and what we mean when we call a function (locally) Lipschitz on this space. For $\varphi\in\mathcal{C}(\operatorname*{\mathcal{E}})$, we define for $u$ in the core and thus by extension on the domain, the energy measure $\Gamma(u)$ which is a Radon measure, see [64, 68], via:
$\int_{X}\varphi\mathop{}\!\mathrm{d}\Gamma(u)\coloneqq\mathcal{E}(\varphi u,u)-\frac{1}{2}\mathcal{E}(u^{2},\varphi).$
If $\Gamma$ is absolutely continuous, this can also be seen as a carré-du-champ-operator if ones looks at the bilinear form obtained using a polarization formula and for $u,v,\varphi\in D(\mathcal{E})$
$\int_{X}\varphi\mathop{}\!\mathrm{d}\Gamma(u,v)=\frac{1}{2}\left(\mathcal{E}(\varphi u,v)+\mathcal{E}(\varphi v,u)-\mathcal{E}(\varphi,uv)\right).$
@@
$d_{\mathcal{E}}(x,y)\coloneqq\sup\left\{u(x)-u(y):u\in C^{0}(M),\frac{\mathop{}\!\mathrm{d}\Gamma(u)}{\mathop{}\!\mathrm{d}\mu}\leq 1\right\}.$
This notion means that $\Gamma(u)$ is absolutely continuous with respect to $\mu$ and that the Radon-Nikodym derivative is bounded by $1$. A priori, this energy metric does not have to be a metric, as it can be infinite or zero for different $x,y\in M$. But for strictly local Dirichlet forms, this is by definition a metric.
We note that the forms $\mathcal{E}_{i}$ are strictly local regular Dirichlet forms, as they are the usual Dirichlet forms on smooth compact manifolds.
[…]
Proposition 2.8 (Characterization Lipschitz functions, [66, Theorem 3.2, 4.1]).
Let $X$ be a locally compact space with strictly local Dirichlet form $\mathcal{E}$. Then $(X,d_{\mathcal{E}})$ is a length space, and the space of Lipschitz functions $\operatorname*{Lip}_{\operatorname*{\mathcal{E}}}$ in the above definition is a sheaf and coincides with the definition through the distance $d_{\mathcal{E}}$.
@@
[…]
For the connectedness of $(X,d,\mu)$ in our sense, we need a Poincaré inequality.
Definition 2.9 ($2$-Poincaré inequality).
We say that $X$ satisfies a $2$-Poincaré inequality if there exists $c>0$ and a dilation factor $\lambda\geq 1$ such that for all integrable functions $f$ and for each ball $B\subseteq X$:
$\fint\limits_{B}|f-f_{B}|^{2}\mathop{}\!\mathrm{d}\mu$ $\leq cr^{2}\fint\limits_{\lambda B}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu.$
Here, $f_{B}$ is the mean of $f$ over $B$:
$f_{B}\coloneqq\fint\limits_{B}f\mathop{}\!\mathrm{d}\mu=\frac{1}{\mu(B)}\int\limits_{B}f\mathop{}\!\mathrm{d}\mu.$
Remark 1.
@@
Note that if $X$ satisfies a Poincaré inequality, it has to be connected up to negligible sets. Assume that it would not be connected up to a null set and let $X_{1}$ be one component with $f=\chi_{X_{1}}$ and $|\nabla f|=0$. Then, this would be a contradiction as $\mu(X_{1})>0$ and $1>f_{B}>0$ for $B$ large enough.
If not specified, we assume the dilation factor $\lambda$ to be one.
[…]
In Section 3, we also show the continuity and ergodicity properties of these objects. For this, we need capacity bounds.
@@
Let $\Omega$ be a (relative) open subset of $X$. We define the relative $(2,\mu)$-capacity on compact sets $K$.
${\text{Cap}}_{2,\mu}(K,\Omega)\coloneqq\inf\left\{\int\limits_{\Omega}|\nabla u|^{2}\mathop{}\!\mathrm{d}\mu:u\in\operatorname*{Lip}_{\operatorname*{\mathcal{E}}}(X)\cap C_{c}(\Omega),u\geq\chi_{K}\right\}.$
This definition then extends to measurable sets by approximation through inner regularity.
Remark 2.
In the case of a glued space, we write ${\text{Cap}}_{2,\mu_{i}}(K,\Omega)$ as a shorthand for ${\text{Cap}}_{2,\mu_{i}}(K\cap M_{i},\Omega\cap M_{i})$.
@@
Additionally, we show that our constructed heat flow connects the entire space. The right notion to make this precise is the notion of ergodicity and irreducibility, see [24]. In the following, let $S_{h}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ be a semigroup.
@@
We say that a measurable set $E\subseteq X$ is $S_{h}$ invariant if for all $f\in L^{2}(X,\mu)$ and $h>0$, we have
$S_{h}(\chi_{E}f)=\chi_{E}S_{h}(f)\ \mu\text{-a.e.}$
Then, $S_{h}$ is called irreducible if for all $S_{h}$ invariant sets $E$, we have $\mu(E)=0$ or $\mu(E^{c})=0$.
@@
An $L^{2}(X,\mu)$ semigroup $S_{t}$ is ergodic if the set function $T_{t}$ defined by
$T_{t}:X\to X,\quad E\mapsto\operatorname*{supp}(S_{t}\chi_{E})$
is ergodic for each fixed $t>0$, that is, if for a measurable set $E$ with $T_{t}^{-1}(E)=E$, we have $\mu(E)=0$ or $\mu(E^{c})=0$.
@@
We say $\omega$ is in the class $A_{2}(M)$ of $2$-Muckenhoupt weights on a smooth compact manifold $M$, if for all balls $B$,
@@
Here, $\lesssim$ mean that the left hand-side is bounded by the right hand-side up to a universal constant.
@@
For the typical weight $\omega=|x|^{\alpha}$ in $\mathbb{R}^{n}$, we have $\omega\in A_{2}$ iff $\alpha\in(-n,n)$.